OpenStudy (anonymous):

Rex mixed x liters of fruit syrup with y liters of water to make a fruit punch. Three times the number of liters of water he mixed was 6 more than twice the number of liters of syrup he mixed. The total number of liters of syrup and water he mixed was 7 liters. Which graph best shows the number of liters of fruit syrup and water that Rex added to make the fruit punch?

OpenStudy (texaschic101):

x + y = 7 --> x = -y + 7 3y = 2x + 6 now sub -y + 7 in for x in the 2nd equation 3y = 2(-y + 7) + 6 3y = -2y + 14 + 6 3y = -2y + 20 3y + 2y = 20 5y = 20 y = 4 now sub 4 in for y in the 1st equation x + y = 7 x + 4 = 7 x = 7 - 4 x = 3 I do not see a graph, however, he added 3 liters of fruit syrup and 4 liters of water.

OpenStudy (anonymous):

Thanks a lot! I was having trouble with this! :)

OpenStudy (texaschic101):

no problem..glad i could help :)